A sum of Rs.2700 is to be given in the form of 63 prizes. If the prize is of either Rs.100 or Rs.25, find the number of prizes of each type.

 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear student


Let the number of prizes of Rs. 100 be =x, and worth rs.25 be y.

Given that,

100x+25y= 2700

Simplify the above equation, we get

4x+y= 108 ….(1)

As per the given question;

x+y= 63 ….(2)

Now, solve equation (1) and (2), we get

3x= 45

x= 15

Now, substitute x= 15 in equation (2), we get

15+y= 63

y=63-15

y=48

Thus, the number of prizes worth Rs. 1oo is 15 and the number of prizes worth Rs. 25 is 48.


Thanks